# Non-planar ends are continuously unforgettable

### v1

@@ 4 Irreducible epimorphisms
So far, we have established that the image of a Dehn twist is a multitwist under the given hypotheses. We now want to show that the image is in fact a Dehn twist. To accomplish this, we will need to use all the hypotheses in the statement of Theorem 4.1, which has not been the case up to this point.
First, we will need a result of the second author characterizing when multitwists are braided, as well as a quick lemma. We say two multitwists $T_{1}$ and $T_{2}$ are braided if they satisfy the braid relation, that is, if $T_{1}T_{2}T_{1}=T_{2}T_{1}T_{2}$. If $a$ and $b$ are two simple closed curves, then $i(a,b)=1$ if and only if $t_{a}$ and $t_{b}$ are braided (see [14, Propositions 3.11 & 3.13]). The following theorem generalizes this fact to multitwists.
@@
If $c$ and $d$ are curves such that $i(c,d)=1$, then there exists $g\in\Gamma_{c}(M)$ such that $g(a)=c$ and $g(b)=d$. Therefore, $\varphi(t_{c})=\varphi(g\,t_{a}\,g^{-1})=\varphi(g\,t_{b}\,g^{-1})=\varphi(t_{d})$. Now, given any two non-separating curves $a^{\prime}$ and $b^{\prime}$, there exist curves $c_{0},c_{1},\ldots,c_{k}$ such that $i(c_{i},c_{i+1})=1$, $c_{0}=a^{\prime}$, and $c_{k}=b^{\prime}$ (see [14, Lemma 4.5]). It follows that $\varphi(t_{c_{i}})=\varphi(t_{c_{i+1}})$, and hence $\varphi(t_{a^{\prime}})=\varphi(t_{b^{\prime}})$. As $\Gamma_{c}(M)$ is generated by non-separating Dehn twists, we have that $\varphi(\Gamma_{c}(M))$ is cyclic. Now, if the genus of $M$ is at least three, then $\Gamma_{c}(M)$ has trivial abelianization—this is an application of the lantern relation, see [14, Section 5.1] for a proof in the finite-type setting that extends to the infinite-type setting verbatim. Now, the image of $\varphi$ is abelian, as it is cyclic, and hence $\varphi$ is trivial. ∎
@@
Let $\mathcal{A}$ be an Alexander chain in $M$ such that $\{t_{a}\}_{a\in\mathcal{A}}$ generates $\Gamma_{c}(M)$, and let $\mathcal{A}^{\prime}=\bigcup_{a\in A}\varphi_{*}(a)$. Note that such a chain exists by Theorem 3.2 and, by Lemma 4.9, the images $\varphi(t_{a})$ are multitwists.
Let $a,b\in\mathcal{A}$, let $a^{\prime}\in\varphi_{*}(a)$, and let $b^{\prime}\in\varphi_{*}(b)$. If $i(a,b)=0$, then $t_{a}$ and $t_{b}$ commute and so do their images $\varphi(t_{a})$ and $\varphi(t_{b})$. Therefore, by Lemma 4.8, $i(a^{\prime},b^{\prime})=0$. We will use the contrapositive of this fact multiple times, so we record it here for reference:
($\star$) if $\text{if }\,i(a^{\prime},b^{\prime})=1\text{ then }\,i(a,b)=1.$
@@
To show that $\mathcal{A}^{\prime}$ is a chain, it is left to show that it is locally finite. First observe that, for $a^{\prime}\in\mathcal{A}^{\prime}$, there are only finitely many curves in $\mathcal{A}^{\prime}$ that have nontrivial geometric intersection with $a^{\prime}$; this follows from the local finiteness of $\mathcal{A}$ together with ($\star$4). Suppose $\mathcal{A}^{\prime}$ fails to be locally finite, so that there exits a curve $c$ and a sequence $\{a_{n}^{\prime}\}_{n\in\mathbb{N}}\subset\mathcal{A}^{\prime}$ such that $i(a_{n}^{\prime},c)>0$ for all $n\in\mathbb{N}$. All but finitely many of the $a_{n}^{\prime}$ have trivial geometric intersection with $a_{1}^{\prime}$, so passing to a subsequence, we may assume that $i(a_{1}^{\prime},a_{n}^{\prime})=0$ for all $n\in\mathbb{N}$. Recursively passing to subsequences, we can assume that $i(a_{n}^{\prime},a_{m}^{\prime})=0$ for all $n,m\in\mathbb{N}$. Let $a_{n}\in\mathcal{A}$ such that $a_{n}^{\prime}\in\varphi_{*}(a)$. Again recursively passing to subsequences as above, we can assume that $i(a_{n},a_{m})=0$ for all $n,m\in\mathbb{N}$. Therefore, as $\{a_{n}\}_{n\in\mathbb{N}}$ is a locally finite multicurve, Lemma 4.7 implies that $\bigcup_{n\in\mathbb{N}}\varphi_{*}(a_{n})$ is a locally finite multicurve. It follows that $\{a_{n}^{\prime}\}_{n\in\mathbb{N}}$ is locally finite as it is contained in $\{\varphi_{*}(a_{n})\}$, implying that only finitely many of the $a_{n}^{\prime}$ have nonzero geometric intersection with $c$, a contradiction. We can conclude that $\mathcal{A}^{\prime}$ is a chain.
We now need to show that $\mathcal{A}^{\prime}$ is filling and tree-like; first, we show it is filling. If there exists a curve or arc $\alpha$ that is disjoint from from every element of $\mathcal{A}^{\prime}$, then $\alpha$ must be fixed by every element of $\varphi(\Gamma_{c}(M))$, as the image of $\Gamma_{c}(M)$ is generated by $\{\varphi(t_{a})\}_{a\in\mathcal{A}}$. By continuity, $\alpha$ must then be fixed by every element in the image of $\varphi$, but $\varphi$ is irreducible, so no such curve or arc exists. Therefore, $\mathcal{A}^{\prime}$ is filling.
Recall that $\mathcal{T}(\mathcal{A}^{\prime})$ is the graph whose vertices are the curves of $\mathcal{A}^{\prime}$ and that two vertices are adjacent if they have geometric intersection one; to show that $\mathcal{A}^{\prime}$ is tree-like, we must show that $\mathcal{T}(\mathcal{A}^{\prime})$ is a tree, i.e., that it is connected and has no cycles. Suppose $a_{1}^{\prime},a_{2}^{\prime},\ldots,a_{k}^{\prime}$ are distinct curves $\mathcal{A}^{\prime}$ forming a cycle in $\mathcal{T}(\mathcal{A}^{\prime})$. Let $a_{1},\ldots,a_{k}\in\mathcal{A}$ such that $a_{i}^{\prime}\in\varphi_{*}(a_{i})$. Reading indices modulo $k$, we have by ($\star$4) that $i(a_{i},a_{i+1})=1$, and using Theorem 4.10, we can conclude that $a_{i}$, $a_{i+1}$, and $a_{i+2}$ are distinct. Therefore, if $\mathcal{T}(\mathcal{A}^{\prime})$ contains a cycle, so does $\mathcal{T}(\mathcal{A})$. It follows that $\mathcal{T}(\mathcal{A}^{\prime})$ cannot have any cycles, as $\mathcal{T}(\mathcal{A})$ is a tree. Finally, $\mathcal{T}(\mathcal{A}^{\prime})$ is connected as $\mathcal{A}^{\prime}$ is filling. This establishes that $\mathcal{T}(\mathcal{A}^{\prime})$ is a tree and that $\mathcal{A}^{\prime}$ is an Alexander chain.
@@
By the construction of $\Sigma_{n}$, the lower genus of $\mathcal{A}_{n}$ agrees with the genus of $\Sigma_{n}$. Moreover, there are pairs $(a_{1},b_{1}),\ldots,(a_{g},b_{g})\in\mathcal{A}_{n}^{2}$ realizing the lower genus of $\mathcal{A}_{n}$ such that any two distinct curves in $\mathcal{A}_{n}\smallsetminus\{a_{1},b_{1},\ldots,a_{g},b_{g}\}$ are disjoint (e.g., take the $a_{i}$ to be the red curves from Figure 2). As the lower genus of $\mathcal{A}_{n}^{\prime}$ is equal to that of $\mathcal{A}_{n}$, the genus of $\Sigma_{n}^{\prime}$ is at least the genus of $\Sigma_{n}$. Now, suppose that the genus of $\Sigma_{n}^{\prime}$ is greater than the genus of $\Sigma_{n}$. Let $a_{i}^{\prime}=\varphi_{*}(a_{i})$ and $b_{i}^{\prime}=\varphi_{*}(b_{i})$, so $(a_{1}^{\prime},b_{1}^{\prime}),\ldots,(a_{g}^{\prime},b_{g}^{\prime})$ is a maximal collection of pairs of curves in $\mathcal{A}_{n}^{\prime}$ realizing its lower genus. Let $F_{i}^{\prime}$ be a closed regular neighborhood of a realization of $\{a_{i}^{\prime},b_{i}^{\prime}\}$, so that $F_{i}^{\prime}$ is a one-hole torus. We can assume that $F_{i}^{\prime}\cap F_{j}^{\prime}=\varnothing$ if $i\neq j$. Then $\Sigma_{n}^{\prime}\smallsetminus\bigcup F_{i}^{\prime}$ has positive genus, implying that there exists a pair of curves in $\mathcal{A}_{n}^{\prime}\smallsetminus\{a_{1}^{\prime},b_{1}^{\prime},\ldots,a_{g}^{\prime},b_{g}^{\prime}\}$ having nontrivial geometric intersection, a contradiction. Therefore, $\Sigma_{n}$ and $\Sigma_{n}^{\prime}$ have the same genus.
@@
Given the setup at hand, for each $n\in\mathbb{N}$, the restriction of $\varphi$ to $\operatorname{MCG}(\Sigma_{n},\partial\Sigma_{n})$ induces a homomorphism $\varphi_{n}\colon\thinspace\operatorname{MCG}(\Sigma_{n},\partial\Sigma_{n})\to\operatorname{MCG}(\Sigma_{n}^{\prime},\partial\Sigma_{n}^{\prime})$. We claim that $\varphi_{n}$ is an isomorphism. As $\Sigma_{n}$ and $\Sigma_{n}^{\prime}$ have the same genus (and it is at least six), we can apply a result of the first author and Souto [6, Theorem 1.1] to see that $\varphi_{n}$ is induced by an embedding $h_{n}\colon\thinspace\Sigma_{n}\to\Sigma_{n}^{\prime}$.
@@
As $\varphi_{m}$ restricts to $\varphi_{n}$ on $\operatorname{MCG}(\Sigma_{n},\partial\Sigma_{n})$, we can choose the embeddings such that $h_{m}$ restricts to $h_{n}$ on $\Sigma_{n}$. We now claim that $h_{n}$ can be chosen to be a homeomorphism. As $\Sigma_{n}^{\prime}$ is filled by curves contained in $h(\Sigma_{n})$, each component of $\Sigma_{n}^{\prime}\smallsetminus h(\Sigma_{n})$ is either a disk or annulus. We claim that each such component is a boundary annulus, i.e., is homotopic to a component of $\partial\Sigma_{n}^{\prime}$. This implies that $h_{n}$ be chosen to be a homeomorphism. Let $m\in\mathbb{N}$ such that $m>n$ and each component of $\Sigma_{m}\smallsetminus\Sigma_{n}$ has positive genus. It follows that each component of the complement of $\Sigma_{m}^{\prime}\smallsetminus h_{m}(\partial\Sigma_{n})$ has positive genus; in particular, if $\delta_{1}$ and $\delta_{2}$ are any two components of $\partial\Sigma_{n}$, then $h_{m}(\delta_{1})$ does not bound a disk and $h_{m}(\delta_{1})$ and $h_{m}(\delta_{2})$ do not co-bound an annulus. The claim now follows, that is, each component of $\Sigma_{n}^{\prime}\smallsetminus h(\Sigma_{n})$ is a boundary annulus, and $h_{n}$ can be chosen to be a homeomorphism. Therefore, $\varphi_{n}\colon\thinspace\operatorname{MCG}(\Sigma_{n},\partial\Sigma_{n})\to\operatorname{MCG}(\Sigma_{n}^{\prime},\partial\Sigma_{n}^{\prime})$ is an isomorphism.
@@ 5 Proof of the main theorem
Let $a$ and $b$ be simple closed curves in $M$ satisfying $i(a,b)=1$. By Lemma 4.9, $\psi(a)$ and $\psi(b)$ are multitwists; let $A$ (resp., $B$) be the multicurve such that $\psi(a)$ (resp., $\psi(b)$) twists about. By Theorem 4.10, either there exist $a^{\prime}\in A$ and $b^{\prime}\in B$ such that $i(a^{\prime},b^{\prime})=1$ or $\psi(a)=\psi(b)$. If the latter case holds, then Lemma 4.11 implies that $\Gamma_{c}(M)<\ker\psi$, which is not the case; hence, the former case holds, implying that $A$ contains a non-separating simple closed curve.
@@
The second equality implies that $g_{1}\in H$, and the last equality guarantees that $g_{1}$ is nontrivial, as $h(a)\neq a$. Let $S_{1}$ be the surface filled by a realization of $\{a,h(a)\}$; note that $S_{1}$ has finite type and contains the support of a representative of $g_{1}$. Thurston’s construction (see [14, Theorem 14.1]) implies that $g_{1}$ is the image of a pseudo-Anosov element under the homomorphism $\operatorname{MCG}(S_{1},\partial S_{1})\to G$ induced by the embedding $S_{1}\hookrightarrow M$ (if $a$ and $h(a)$ are disjoint, then $S_{1}$ is a union of disjoint annuli implying that $g_{1}$ is a multitwist, which we will also consider pseudo-Anosov in this disjoint union of annuli). It follows that if a curve $c$ has nontrivial geometric intersection with $\partial S_{1}$, then $g_{1}(c)\neq c$; in particular, $g_{1}(c)=c$ if and only if $i(c,\partial S)=0$.

### v2

@@ 4 Irreducible epimorphisms
So far, we have established that the image of a Dehn twist is a multitwist under the given hypotheses. We now want to show that the image is in fact a Dehn twist. To accomplish this, we will need to use irreducibility and the non-planarity of the ends, which has not been the case up to this point.
First, we will need a result of the second author characterizing when multitwists are braided, as well as a quick lemma. We say two multitwists $T_{1}$ and $T_{2}$ are braided if they satisfy the braid relation, that is, if $T_{1}T_{2}T_{1}=T_{2}T_{1}T_{2}$. If $a$ and $b$ are two simple closed curves, then $i(a,b)=1$ if and only if $t_{a}$ and $t_{b}$ are braided (see [15, Propositions 3.11 & 3.13]). The following theorem generalizes this fact to multitwists.
@@
If $c$ and $d$ are curves such that $i(c,d)=1$, then there exists $g\in\Gamma_{c}(M)$ such that $g(a)=c$ and $g(b)=d$. Therefore, $\varphi(t_{c})=\varphi(g\,t_{a}\,g^{-1})=\varphi(g\,t_{b}\,g^{-1})=\varphi(t_{d})$. Now, given any two non-separating curves $a^{\prime}$ and $b^{\prime}$, there exist curves $c_{0},c_{1},\ldots,c_{k}$ such that $i(c_{i},c_{i+1})=1$, $c_{0}=a^{\prime}$, and $c_{k}=b^{\prime}$ (see [15, Lemma 4.5]). It follows that $\varphi(t_{c_{i}})=\varphi(t_{c_{i+1}})$, and hence $\varphi(t_{a^{\prime}})=\varphi(t_{b^{\prime}})$. As $\Gamma_{c}(M)$ is generated by non-separating Dehn twists, we have that $\varphi(\Gamma_{c}(M))$ is cyclic. Now, if the genus of $M$ is at least three, then $\Gamma_{c}(M)$ has trivial abelianization—this is an application of the lantern relation, see [15, Section 5.1] for a proof in the finite-type setting that extends to the infinite-type setting verbatim. Now, the image of $\varphi$ is abelian, as it is cyclic, and hence $\varphi$ is trivial. ∎
@@
Apply Theorem 3.2 to obtain an Alexander chain $\mathcal{A}$ in $M$ such that $\{t_{a}\}_{a\in\mathcal{A}}$ generates $\Gamma_{c}(M)$, and let $\mathcal{A}^{\prime}=\bigcup_{a\in A}\varphi_{*}(a)$. Note that, by Lemma 4.9, the images $\varphi(t_{a})$ of the $t_{a}$ are multitwists.
Let $a,b\in\mathcal{A}$, let $a^{\prime}\in\varphi_{*}(a)$, and let $b^{\prime}\in\varphi_{*}(b)$. If $i(a,b)=0$, then $t_{a}$ and $t_{b}$ commute, and $\varphi(t_{a})$ and $\varphi(t_{b})$ are commuting multitwists. Therefore, by Lemma 2.5, $i(a^{\prime},b^{\prime})=0$. We will use the contrapositive of this fact multiple times, so we record it here for reference:
($\star$) $\text{if }\,i(a^{\prime},b^{\prime})\neq 0\text{ then }\,i(a,b)=1.$
@@
To show that $\mathcal{A}^{\prime}$ is a chain, it is left to show that it is locally finite. As $\mathcal{A}$ is given by the construction in Theorem 3.2, we can write $\mathcal{A}=A_{1}\cup A_{2}\cup A_{3}$, where $A_{1}$, $A_{2}$, and $A_{3}$ are pairwise disjoint locally finite multicurves (e.g., take $A_{1}$ to be the red curves, $A_{2}$ to be the blue curves, and $A_{3}$ to be the remaining curves, which are purple in Figure 2). By Lemma 4.7, $\varphi_{*}(A_{i})$ is locally finite; hence, $\mathcal{A}^{\prime}=\varphi_{*}(A_{1})\cup\varphi_{*}(A_{2})\cup\varphi_{*}(A_{3})$ expresses $\mathcal{A}^{\prime}$ as a finite union of locally finite sets, implying $\mathcal{A}^{\prime}$ is itself locally finite. Therefore, $\mathcal{A}^{\prime}$ is a chain.
Recall that $\mathcal{T}(\mathcal{A}^{\prime})$ is the graph whose vertices are the curves of $\mathcal{A}^{\prime}$ and that two vertices are adjacent if they have geometric intersection one; to show that $\mathcal{A}^{\prime}$ is tree-like, we must show that $\mathcal{T}(\mathcal{A}^{\prime})$ is a tree, i.e., that it is connected and has no cycles. Suppose $a_{1}^{\prime},a_{2}^{\prime},\ldots,a_{k}^{\prime}$ are distinct curves $\mathcal{A}^{\prime}$ forming a cycle in $\mathcal{T}(\mathcal{A}^{\prime})$. Let $a_{1},\ldots,a_{k}\in\mathcal{A}$ such that $a_{i}^{\prime}\in\varphi_{*}(a_{i})$. Reading indices modulo $k$, we have by ($\star$4) that $i(a_{i},a_{i+1})=1$. We claim that $a_{i}$, $a_{i+1}$, and $a_{i+2}$ are distinct. To see this, observe that as $\varphi(t_{a_{i}})$ and $\varphi(t_{a_{i+1}})$ are braided, Theorem 4.10 implies that each component of $\varphi_{*}(t_{a_{i+1}})$ has nontrivial geometric intersection with exactly one component of $\varphi_{*}(t_{a_{i}})$. In particular, as $i(a_{i}^{\prime},a_{i+1}^{\prime})=1$ and $i(a_{i+1}^{\prime},a_{i+2}^{\prime})=1$, we can conclude that $a_{i+2}^{\prime}\notin\varphi_{*}(a_{i})$, allowing us to conclude that $a_{i}$, $a_{i+1}$, and $a_{i+2}$ are distinct curves. Therefore, if $\mathcal{T}(\mathcal{A}^{\prime})$ contains a cycle, so does $\mathcal{T}(\mathcal{A})$. It follows that $\mathcal{T}(\mathcal{A}^{\prime})$ cannot have any cycles, as $\mathcal{T}(\mathcal{A})$ is a tree. Finally, $\mathcal{T}(\mathcal{A}^{\prime})$ is connected as $\mathcal{A}^{\prime}$ is filling. This establishes that $\mathcal{T}(\mathcal{A}^{\prime})$ is a tree and that $\mathcal{A}^{\prime}$ is an Alexander chain.
@@
By the construction of $\Sigma_{n}$, the lower genus of $\mathcal{A}_{n}$ agrees with the genus of $\Sigma_{n}$. Moreover, there are pairs $(a_{1},b_{1}),\ldots,(a_{g},b_{g})\in\mathcal{A}_{n}^{2}$ realizing the lower genus of $\mathcal{A}_{n}$ such that any two distinct curves in $\mathcal{A}_{n}\smallsetminus\{a_{1},b_{1},\ldots,a_{g},b_{g}\}$ are disjoint (e.g., take the $a_{i}$ to be the red curves from Figure 3). As the lower genus of $\mathcal{A}_{n}^{\prime}$ is equal to that of $\mathcal{A}_{n}$, the genus of $\Sigma_{n}^{\prime}$ is at least the genus of $\Sigma_{n}$. Now, suppose that the genus of $\Sigma_{n}^{\prime}$ is greater than the genus of $\Sigma_{n}$. Let $a_{i}^{\prime}=\varphi_{*}(a_{i})$ and $b_{i}^{\prime}=\varphi_{*}(b_{i})$, so $(a_{1}^{\prime},b_{1}^{\prime}),\ldots,(a_{g}^{\prime},b_{g}^{\prime})$ is a maximal collection of pairs of curves in $\mathcal{A}_{n}^{\prime}$ realizing its lower genus. Let $F_{i}^{\prime}$ be a closed regular neighborhood of a realization of $\{a_{i}^{\prime},b_{i}^{\prime}\}$, so that $F_{i}^{\prime}$ is a one-holed torus. We can assume that $F_{i}^{\prime}\cap F_{j}^{\prime}=\varnothing$ if $i\neq j$. Then $\Sigma_{n}^{\prime}\smallsetminus\bigcup F_{i}^{\prime}$ has positive genus, implying that there exists a pair of curves in $\mathcal{A}_{n}^{\prime}\smallsetminus\{a_{1}^{\prime},b_{1}^{\prime},\ldots,a_{g}^{\prime},b_{g}^{\prime}\}$ having nontrivial geometric intersection, a contradiction. Therefore, $\Sigma_{n}$ and $\Sigma_{n}^{\prime}$ have the same genus.
@@
Given the setup at hand, for each $n\in\mathbb{N}$, the restriction of $\varphi$ to $\operatorname{MCG}(\Sigma_{n},\partial\Sigma_{n})$ induces a homomorphism $\varphi_{n}\colon\thinspace\operatorname{MCG}(\Sigma_{n},\partial\Sigma_{n})\to\operatorname{MCG}(\Sigma_{n}^{\prime},\partial\Sigma_{n}^{\prime})$. We claim that $\varphi_{n}$ is an . As $\Sigma_{n}$ and $\Sigma_{n}^{\prime}$ have the same genus (and it is at least six), we can apply a result of the first author and Souto [6, Theorem 1.1] to see that $\varphi_{n}$ is induced by an embedding $h_{n}\colon\thinspace\Sigma_{n}\to\Sigma_{n}^{\prime}$.
@@
As $\varphi_{m}$ restricts to $\varphi_{n}$ on $\operatorname{MCG}(\Sigma_{n},\partial\Sigma_{n})$ for $m>n$, we can choose the embeddings such that $h_{m}$ restricts to $h_{n}$ on $\Sigma_{n}$. We now claim that $h_{n}$ can be chosen to be a homeomorphism. As $\Sigma_{n}^{\prime}$ is filled by curves contained in $h_{n}(\Sigma_{n})$, each component of $\Sigma_{n}^{\prime}\smallsetminus h_{n}(\Sigma_{n})$ is either a disk or annulus. We claim that each such component is a boundary annulus, i.e., is homotopic to a component of $\partial\Sigma_{n}^{\prime}$. This implies that $h_{n}$ can be chosen to be a homeomorphism. Let $m\in\mathbb{N}$ such that $m>n$ and each component of $\Sigma_{m}\smallsetminus\Sigma_{n}$ has positive genus. It follows that each component of the complement of $\Sigma_{m}^{\prime}\smallsetminus h_{m}(\partial\Sigma_{n})$ has positive genus; in particular, if $\delta_{1}$ and $\delta_{2}$ are any two components of $\partial\Sigma_{n}$, then $h_{m}(\delta_{1})$ does not bound a disk and $h_{m}(\delta_{1})$ and $h_{m}(\delta_{2})$ do not co-bound an annulus. The claim now follows, that is, each component of $\Sigma_{n}^{\prime}\smallsetminus h_{n}(\Sigma_{n})$ is a boundary annulus, and $h_{n}$ can be chosen to be a homeomorphism. Therefore, $\varphi_{n}\colon\thinspace\operatorname{MCG}(\Sigma_{n},\partial\Sigma_{n})\to\operatorname{MCG}(\Sigma_{n}^{\prime},\partial\Sigma_{n}^{\prime})$ is an isomorphism.
@@ 5 Proof of the main theorem
Let $a$ and $b$ be simple closed curves in $M$ satisfying $i(a,b)=1$. By Lemma 4.9, $\psi(t_{a})$ and $\psi(t_{b})$ are multitwists; let $A$ (resp., $B$) be the multicurve such that $\psi(t_{a})$ (resp., $\psi(t_{b})$) twists about. Note that $\psi(t_{a})$ and $\psi(t_{b})$ are braided multitwists, and hence by Theorem 4.10, either there exist $a^{\prime}\in A$ and $b^{\prime}\in B$ such that $i(a^{\prime},b^{\prime})=1$ or $\psi(t_{a})=\psi(t_{b})$. If the latter case holds, then Lemma 4.11 implies that $\Gamma_{c}(M)<\ker\psi$, which is not the case; hence, the former case holds, implying that $A$ contains a non-separating simple closed curve.
@@
The second equality implies that $g_{1}\in H$, and the last equality guarantees that $g_{1}$ is nontrivial, as $h(a)\neq a$. Let $S_{1}$ be the surface filled by a realization of $\{a,h(a)\}$; note that $S_{1}$ has finite type and contains the support of a representative of $g_{1}$. Thurston’s construction (see [15, Theorem 14.1]) implies that $g_{1}$ is the image of a pseudo-Anosov element under the homomorphism $\operatorname{MCG}(S_{1},\partial S_{1})\to G$ induced by the embedding $S_{1}\hookrightarrow M$ (if $a$ and $h(a)$ are disjoint, then $S_{1}$ is a union of disjoint annuli implying that $g_{1}$ is a multitwist, which we will also consider pseudo-Anosov in this disjoint union of annuli). It follows that if a curve $c$ has nontrivial geometric intersection with $\partial S_{1}$, then $g_{1}(c)\neq c$; in particular, $g_{1}(c)=c$ if and only if $i(c,\partial S)=0$.